# CoD: A Diffusion Foundation Model for Image Compression

### v2

@@ 5.1 Generative Image Compression
Unlike neural image codecs [balle2017end, hyperprior, elic] that primarily minimize reconstruction distortion (MSE), generative image compression [agustsson2019generative, iwai2021fidelity, po_elic, glc, multi_realism, dlf] aims to achieve both low distortion and high perceptual quality. Early works [hific, illm] introduced conditional GANs to enhance perceptual realism, successfully improving perceptual quality compared to MSE-optimized codecs, albeit at the cost of higher reconstruction distortion. TACO [taco] further incorporates textual information to mitigate MSE degradation, yet still falls short in achieving an ideal distortion–perception balance. To the best of our knowledge, we are the first to demonstrate that pixel-space diffusion can attain VTM-level distortion while delivering superior perceptual quality to GAN-based codecs.
@@ 5.2 Diffusion Image Compression
Recent advances in generative image compression increasingly leverage the powerful generative priors of diffusion models to achieve higher realism. PerCo [perco] compresses image representations and text captions as conditions, fine-tuning a pretrained latent diffusion model for perfect realism. In contrast, DiffEIC [diffeic] and OSCAR [oscar] show that removing the text bitstream can yield better overall performance. Other approaches [psc, idempotence] investigate zero-shot compression using pretrained diffusion models without fine-tuning. DiffC [diffc_original, diffc] introduces reverse-channel coding to progressively compress increasingly noised representations, while DDCM [ddcm] adopts random codebooks to encode the random sampled noises. Most of these methods depend on pretrained diffusion foundation models such as Stable Diffusion, which are theoretically suboptimal for compression. CoD offers an alternative, compression-oriented diffusion foundation model that can serve as a stronger backbone for such approaches, enabling superior performance.
Another early exploration on pixel-space diffusion codec is CDC [cdc], which also formulates compression as conditions for diffusion. However, CoD differs in several key aspects. First, CoD emphasizes training at ultra-low bitrates ($\sim$0.0039 bpp) to strengthen its strong generative prior, and it relies purely on diffusion loss rather than traditional perceptual losses. In contrast, CDC is trained at high bitrates (>>0.2 bpp), where dense information limits generative learning. Consequently, CDC requires perceptual losses such as LPIPS, otherwise its perceptual quality drops to the level of an MSE-optimized codec. Second, CoD embodies the concept of a foundation model, scaling training to push the frontier of compression research, while CDC does not consider scaling laws. Finally, CoD offers insights beyond the model itself: we provide a comprehensive analysis of scaling behavior, pixel- vs. latent-space and conditioning modalities, offering broader insights for future research.
@@ 6 Conclusion and Limitation
  • One-step diffusion codec: Finetuning one-step CoD achieves real-time and SOTA performance, comparable to StableCodec [stablecodec] and OneDC [onedc].
  • CoD as perceptual loss: CoD-based perceptual loss significantly improves MS-ILLM [illm].
@@ Appendix A Towards 64-Bit Image Compression
We evaluate our 64-bit CoD on the Kodak dataset in Figure 10. Surprisingly, under such an extremely low bitrate, CoD successfully reconstructs the correct semantics of the original image. By contrast, Stable Diffusion based codecs DDCM [ddcm] and DiffC [diffc] fail to reconstruct correct semantics even at a fourfold ($4\times$) higher bit cost. We further leverage DiffC to evaluate the downstream compression performance on Kodak [kodak] in Figure 11, where CoD-based DiffC significantly outperforms other Stable Diffusion based codecs. Our scheme achieves a FID of 70 with only less than $10\%$ of the bits of prior codecs, highlighting the extreme compression capability of 64-bit CoD.
@@
Figure 14: Rate-Distortion curves for MS-ILLM [illm] and finetuning MS-ILLM decoder with CoD-based DMD loss on Kodak at $512\times 512$.
@@ B.1 Preliminary
Recently, one-step diffusion models have demonstrated effectiveness in generative image compression. Unlike multi-step diffusion, which optimizes each diffusion timestep separately, one-step diffusion can be trained end-to-end, enabling better overall performance. OSCAR [oscar] fine-tuned a multi-step Stable Diffusion to function as a one-step image codec. StableCodec [stablecodec] leverages a one-step SD-Turbo as the foundation model, achieving significantly improved performance. Similarly, OneDC [onedc] employs a one-step DMD-distilled [dmd] Stable Diffusion as the foundation model for one-step diffusion compression. The success of these approaches motivates our exploration of CoD in the context of one-step diffusion foundation models.
Since the SD-Turbo distillation process is not fully public, we follow DMD [dmd] to distill our diffusion model. Given a one-step diffusion network $G_{\theta}(z)$ that generates $x$ using random noise $z$, DMD proves that the KL divergence between the real and fake distributions can be expressed as a distribution-matching loss:
@@
where $s_{\mathrm{real}}(x)$ and $s_{\mathrm{fake}}(x)$ are the score functions [score_matching] of their respective distributions. By adding random noise to $x$, the score can be predicted by a diffusion model. DMD leverages a pretrained multi-step diffusion model for the real score and dynamically trains a diffusion model to estimate the fake score for one-step samples $x$.
@@ B.2 Distilling a CoD
where $L_{\mathrm{pixel}}=\mathcal{L}_{1}+\mathcal{L}_{\mathrm{LPIPS}}$ consists of pixel-space L1 loss and the LPIPS [lpips] loss, $L_{\mathrm{perc}}=2\,\cdot\,\mathcal{L}_{\mathrm{DMD}}+0.01\,\cdot\,\mathcal{L}_{\mathrm{GAN}}$ consists of a DMD loss and a patch GAN loss [vqgan], $\mathcal{L}_{\mathrm{REPA}}$ is the representation alignment loss and $\mathcal{L}_{\mathrm{C}}$ is th codebook commitment loss. We set $\lambda=0.5$ and $\beta=0.25$.
@@ B.3 Finetuning to Higher Bitrates
Stage I: initialization. We randomly set the parameters of the condition encoder, codebook, and condition decoder, while loading the pre-trained multi-step CoD weights for the diffusion module. During this warm-up phase, the diffusion module is trained using LoRA [lora] with rank 32. To accelerate convergence, we remove the $L_{\mathrm{perc}}$ term from $L_{\mathrm{OS}}$ and train for 200K steps with a learning rate of $10^{-4}$.
@@ B.4 Evaluation
Comparison with one-step codecs. We evaluate the fine-tuned one-step CoD on the Kodak dataset (Figure 12). It yields competitive reconstruction quality compared with SOTA one-step diffusion codecs such as OneDC and StableCodec, particularly at ultra-low bitrates. Notably, our method does not rely on any pretrained components from Stable Diffusion (e.g., one-step models or SD-VAEs) to reach this performance. We currently adopt fixed-length coding for the codebook indices for simplicity. Further gains can be achieved by introducing an entropy model or applying latent compression [glc]. Visual comparisons in Figure 13 further show that one-step CoD preserves higher fidelity than both multi-step CoD and OneDC.
@@ Appendix C Towards Perceptual Supervision via CoD
In the DMD loss in Equation 9, for any input $x$, we can optimize its KL divergence using CoD to estimate real and fake scores. This indicates that CoD can act as a perceptual supervision mechanism, enhancing the realism of reconstructions for a variety of networks, including other pixel-space codecs, restoration models, or super-resolution models. To demonstrate this potential, we finetune the pixel-space codec MS-ILLM [illm] using the CoD-based DMD loss.
@@
Evaluation. MS-ILLM is typically pretrained using the MSE loss, after which the encoder and entropy model are fixed while the decoder is finetuned with perceptual losses such as LPIPS and GAN [gan] loss. Following this paradigm, we finetune only the decoder of MS-ILLM, replacing the GAN loss with the CoD-based DMD loss. Remarkably, after optimizing the decoder for just 2K steps with a batch size of 32 (20K steps total including fake score learning), MS-ILLM demonstrates substantially improved perceptual quality across all metrics, as shown in Figure 14. Towards ultra-low bitrates like 0.011 bpp where the information is severely distorted, we find increasing the learning rate of fake score network training can better capture the distorted distribution to enhance realism, as shown in Table 1. Several visual examples are presented in Figure 15, where the CoD-based DMD loss markedly reduces artifacts and yields clearer edges and finer details. Since the encoder is fixed, the encoded information remains heavily distorted under MSE-only optimization, resulting in poor overall reconstruction. We expect that jointly tuning the entire network with the CoD-based DMD loss could further improve perceptual quality.
@@ D.2 Finetuning Multi-step CoD at Higher Bitrates
A straightforward downstream compression scheme is to directly finetune multi-step CoD at higher bitrates through diffusion loss. However, this leads to inferior results compared to other schemes. When jointly optimizing a condition under explicit diffusion loss, the optimization becomes biased: the condition tends to maximize the score norm, as shown in [max_score_norm]. In practice, this drives the decoder toward oversharpened and over-saturated reconstructions, resulting in lower fidelity. This observation is consistent with recent findings that one-step diffusion codecs perform better because they avoid this biased optimization.
@@ Appendix E Exploring $\mathcal{X}$-Prediction Pixel Diffusion
Since $\mathcal{V}$-prediction (velocity prediction) was proposed [v-pred] and later adopted by Stable Diffusion v2.1, it has become a common practice in latent diffusion models, including the baseline diffusion architectures used in our work [ddt, pixnerd]. However, recent research by Just Image Diffusion (JiT) [jit] shows that pixels lie on a low-dimensional manifold, making $\mathcal{X}$-prediction a more effective approach for directly modeling pixel distributions. In this paper, we primarily adopt $\mathcal{V}$-prediction to provide a unified view of latent- and pixel-space compression. Nevertheless, in this section, we conduct experiments to demonstrate that $\mathcal{X}$-prediction can achieve better reconstruction performance.
@@
Comparison on pixel-space CoD. We follow the JiT training pipeline to build a pixel-space CoD variant using $\mathcal{X}$-prediction and $\mathcal{V}$-loss, denoted as CoD (pixel-$\mathcal{X}$) to distinguish it from our velocity-based pixel-space CoD. To further examine the influence of network design (i.e., the decoupled head [ddt, pixnerd]) we also substitute our pixel-space diffusion model with JiT’s DiT backbone, referred to as CoD (pixel-$\mathcal{X}$, JiT). As shown in Table 2, $\mathcal{X}$-prediction yields significantly better perceptual metrics than $\mathcal{V}$-prediction. In contrast, replacing the decoupled-head design with JiT’s pure DiT structure does not provide performance gains.
@@
Appendix F Additional Details and Results
This section provides further training details and additional evaluation results, including complexity analysis and visual comparisons.
F.1 Training Details
F.1.1 Datasets
We train the model using three publicly available datasets: ImageNet-21K [imagenet] contains 14.2M images across 21K categories. After removing images with a shorter edge below 256 pixels, 9.3M images remain for $256\times 256$ pre-training. OpenImages [openimage] and SA-1B [sa1b] provide high-resolution images. We use 1.7M directly downloadable images from OpenImages V4 and all 11.1M images from SA-1B for both $256\times 256$ and $512\times 512$ resolution pre-training. Overall, CoD is trained on 22.1M images. Compared with modern diffusion pipelines, this is a relatively modest data scale, and the only filtering criterion is image resolution rather than extensive data cleaning. We expect that scaling up training with higher-quality data will further improve the performance of CoD.
F.1.2 Unified Training
ID Ablation @ 0.0039 bpp PSNR LPIPS FID
A Flow matching loss 9.83 dB 0.576 76.8
B A + Auxiliary Loss 15.20 dB 0.458 48.3
C A + Unified Training 15.83 dB 0.433 48.4
D B + Unified Post-Training 16.20 dB 0.433 46.0
@@ F.1.2 Unified Training
    Using BLIP2 captioner with at most 32 tokens, as suggested by PerCo (SD) [perco_sd].
@@
Auxiliary Loss. Given the output of the condition decoder, we attach two lightweight auxiliary prediction heads to (1) reconstruct the input $x$ and (2) predict its DINO v2 [dino_v2] representation. The auxiliary objective is
@@
Unified Post-Training. Motivated by the complementary strengths of unified training and auxiliary supervision, we combine both approaches to further improve performance. Specifically, we first pre-train CoD with auxiliary loss and then post-train it using unified training. As shown in Table 3, this unified post-training strategy (ID = D) achieves the best overall results among all configurations.

### v3

@@ 5.1 Generative Image Compression
Unlike neural image codecs [4, 5, 16] that primarily minimize reconstruction distortion (MSE), generative image compression [3, 23, 17, 24, 1, 52] aims to achieve both low distortion and high perceptual quality. Early works [35, 36] introduced conditional GANs to enhance perceptual realism, successfully improving perceptual quality compared to MSE-optimized codecs, albeit at the cost of higher reconstruction distortion. TACO [28] further incorporates textual information to mitigate MSE degradation, yet still falls short in achieving an ideal distortion–perception balance. To the best of our knowledge, we are the first to demonstrate that pixel-space diffusion can attain VTM-level distortion while delivering superior perceptual quality to GAN-based codecs.
@@ 5.2 Diffusion Image Compression
Recent advances in generative image compression increasingly leverage the powerful generative priors of diffusion models to achieve higher realism. PerCo [7] compresses image representations and text captions as conditions, fine-tuning a pretrained latent diffusion model for perfect realism. In contrast, DiffEIC [32] and OSCAR [15] show that removing the text bitstream can yield better overall performance. Other approaches [11, 50] investigate zero-shot compression using pretrained diffusion models without fine-tuning. DiffC [43, 46] introduces reverse-channel coding to progressively compress increasingly noised representations, while DDCM [37] adopts random codebooks to encode the random sampled noises. Most of these methods depend on pretrained diffusion foundation models such as Stable Diffusion, which are theoretically suboptimal for compression. CoD offers an alternative, compression-oriented diffusion foundation model that can serve as a stronger backbone for such approaches, enabling superior performance.
Another early exploration on pixel-space diffusion codec is CDC [54], which also formulates compression as conditions for diffusion. However, CoD differs in several key aspects. First, CoD emphasizes training at ultra-low bitrates ($\sim$0.0039 bpp) to strengthen its strong generative prior, and it relies purely on diffusion loss rather than traditional perceptual losses. In contrast, CDC is trained at high bitrates (>>0.2 bpp), where dense information limits generative learning. Consequently, CDC requires perceptual losses such as LPIPS, otherwise its perceptual quality drops to the level of an MSE-optimized codec. Second, CoD embodies the concept of a foundation model, scaling training to push the frontier of compression research, while CDC does not consider scaling laws. Finally, CoD offers insights beyond the model itself: we provide a comprehensive analysis of scaling behavior, pixel- vs. latent-space and conditioning modalities, offering broader insights for future research.
@@ 6 Conclusion and Limitation
  • v3 (2026-03-16): Updated to CVPR camera ready version. Updated Github code link.
@@
  • One-step diffusion codec: Finetuning one-step CoD achieves real-time and SOTA performance, comparable to StableCodec [58] and OneDC [53].
  • CoD as perceptual loss: CoD-based perceptual loss significantly improves MS-ILLM [36].
@@ Appendix A Towards 64-Bit Image Compression
We evaluate our 64-bit CoD on the Kodak dataset in Figure 10. Surprisingly, under such an extremely low bitrate, CoD successfully reconstructs the correct semantics of the original image. By contrast, Stable Diffusion based codecs DDCM [37] and DiffC [46] fail to reconstruct correct semantics even at a fourfold ($4\times$) higher bit cost. We further leverage DiffC to evaluate the downstream compression performance on Kodak [10] in Figure 11, where CoD-based DiffC significantly outperforms other Stable Diffusion based codecs. Our scheme achieves a FID of 70 with only less than $10\%$ of the bits of prior codecs, highlighting the extreme compression capability of 64-bit CoD.
@@
Figure 14: Rate-Distortion curves for MS-ILLM [36] and finetuning MS-ILLM decoder with CoD-based DMD loss on Kodak at $512\times 512$.
@@ B.1 Preliminary
Recently, one-step diffusion models have demonstrated effectiveness in generative image compression. Unlike multi-step diffusion, which optimizes each diffusion timestep separately, one-step diffusion can be trained end-to-end, enabling better overall performance. OSCAR [15] fine-tuned a multi-step Stable Diffusion to function as a one-step image codec. StableCodec [58] leverages a one-step SD-Turbo as the foundation model, achieving significantly improved performance. Similarly, OneDC [53] employs a one-step DMD-distilled [55] Stable Diffusion as the foundation model for one-step diffusion compression. The success of these approaches motivates our exploration of CoD in the context of one-step diffusion foundation models.
Since the SD-Turbo distillation process is not fully public, we follow DMD [55] to distill our diffusion model. Given a one-step diffusion network $G_{\theta}(z)$ that generates $x$ using random noise $z$, DMD proves that the KL divergence between the real and fake distributions can be expressed as a distribution-matching loss:
@@
where $s_{\mathrm{real}}(x)$ and $s_{\mathrm{fake}}(x)$ are the score functions [42] of their respective distributions. By adding random noise to $x$, the score can be predicted by a diffusion model. DMD leverages a pretrained multi-step diffusion model for the real score and dynamically trains a diffusion model to estimate the fake score for one-step samples $x$.
@@ B.2 Distilling a CoD
where $L_{\mathrm{pixel}}=\mathcal{L}_{1}+\mathcal{L}_{\mathrm{LPIPS}}$ consists of pixel-space L1 loss and the LPIPS [57] loss, $L_{\mathrm{perc}}=2\,\cdot\,\mathcal{L}_{\mathrm{DMD}}+0.01\,\cdot\,\mathcal{L}_{\mathrm{GAN}}$ consists of a DMD loss and a patch GAN loss [12], $\mathcal{L}_{\mathrm{REPA}}$ is the representation alignment loss and $\mathcal{L}_{\mathrm{C}}$ is th codebook commitment loss. We set $\lambda=0.5$ and $\beta=0.25$.
@@ B.3 Finetuning to Higher Bitrates
Stage I: initialization. We randomly set the parameters of the condition encoder, codebook, and condition decoder, while loading the pre-trained multi-step CoD weights for the diffusion module. During this warm-up phase, the diffusion module is trained using LoRA [22] with rank 32. To accelerate convergence, we remove the $L_{\mathrm{perc}}$ term from $L_{\mathrm{OS}}$ and train for 200K steps with a learning rate of $10^{-4}$.
@@ B.4 Evaluation
Comparison with one-step codecs. We evaluate the fine-tuned one-step CoD on the Kodak dataset (Figure 12). It yields competitive reconstruction quality compared with SOTA one-step diffusion codecs such as OneDC and StableCodec, particularly at ultra-low bitrates. Notably, our method does not rely on any pretrained components from Stable Diffusion (e.g., one-step models or SD-VAEs) to reach this performance. We currently adopt fixed-length coding for the codebook indices for simplicity. Further gains can be achieved by introducing an entropy model or applying latent compression [24]. Visual comparisons in Figure 13 further show that one-step CoD preserves higher fidelity than both multi-step CoD and OneDC.
@@ Appendix C Towards Perceptual Supervision via CoD
In the DMD loss in Equation 9, for any input $x$, we can optimize its KL divergence using CoD to estimate real and fake scores. This indicates that CoD can act as a perceptual supervision mechanism, enhancing the realism of reconstructions for a variety of networks, including other pixel-space codecs, restoration models, or super-resolution models. To demonstrate this potential, we finetune the pixel-space codec MS-ILLM [36] using the CoD-based DMD loss.
@@
Evaluation. MS-ILLM is typically pretrained using the MSE loss, after which the encoder and entropy model are fixed while the decoder is finetuned with perceptual losses such as LPIPS and GAN [14] loss. Following this paradigm, we finetune only the decoder of MS-ILLM, replacing the GAN loss with the CoD-based DMD loss. Remarkably, after optimizing the decoder for just 2K steps with a batch size of 32 (20K steps total including fake score learning), MS-ILLM demonstrates substantially improved perceptual quality across all metrics, as shown in Figure 14. Towards ultra-low bitrates like 0.011 bpp where the information is severely distorted, we find increasing the learning rate of fake score network training can better capture the distorted distribution to enhance realism, as shown in Table 1. Several visual examples are presented in Figure 15, where the CoD-based DMD loss markedly reduces artifacts and yields clearer edges and finer details. Since the encoder is fixed, the encoded information remains heavily distorted under MSE-only optimization, resulting in poor overall reconstruction. We expect that jointly tuning the entire network with the CoD-based DMD loss could further improve perceptual quality.
@@ D.2 Finetuning Multi-step CoD at Higher Bitrates
A straightforward downstream compression scheme is to directly finetune multi-step CoD at higher bitrates through diffusion loss. However, this leads to inferior results compared to other schemes. When jointly optimizing a condition under explicit diffusion loss, the optimization becomes biased: the condition tends to maximize the score norm, as shown in [51]. In practice, this drives the decoder toward oversharpened and over-saturated reconstructions, resulting in lower fidelity. This observation is consistent with recent findings that one-step diffusion codecs perform better because they avoid this biased optimization.
@@ Appendix E Exploring $\mathcal{X}$-Prediction Pixel Diffusion
Since $\mathcal{V}$-prediction (velocity prediction) was proposed [41] and later adopted by Stable Diffusion v2.1, it has become a common practice in latent diffusion models, including the baseline diffusion architectures used in our work [49, 48]. However, recent research by Just Image Diffusion (JiT) [31] shows that pixels lie on a low-dimensional manifold, making $\mathcal{X}$-prediction a more effective approach for directly modeling pixel distributions. In this paper, we primarily adopt $\mathcal{V}$-prediction to provide a unified view of latent- and pixel-space compression. Nevertheless, in this section, we conduct experiments to demonstrate that $\mathcal{X}$-prediction can achieve better reconstruction performance.
@@
Comparison on pixel-space CoD. We follow the JiT training pipeline to build a pixel-space CoD variant using $\mathcal{X}$-prediction and $\mathcal{V}$-loss, denoted as CoD (pixel-$\mathcal{X}$) to distinguish it from our velocity-based pixel-space CoD. To further examine the influence of network design (i.e., the decoupled head [49, 48]) we also substitute our pixel-space diffusion model with JiT’s DiT backbone, referred to as CoD (pixel-$\mathcal{X}$, JiT). As shown in Table 2, $\mathcal{X}$-prediction yields significantly better perceptual metrics than $\mathcal{V}$-prediction. In contrast, replacing the decoupled-head design with JiT’s pure DiT structure does not provide performance gains.
@@
    Using BLIP2 captioner with at most 32 tokens, as suggested by PerCo (SD) [26].
@@
ID Ablation @ 0.0039 bpp PSNR LPIPS FID
A Flow matching loss 9.83 dB 0.576 76.8
B A + Auxiliary Loss 15.20 dB 0.458 48.3
C A + Unified Training 15.83 dB 0.433 48.4
D B + Unified Post-Training 16.20 dB 0.433 46.0
Appendix F Additional Details and Results
This section provides further training details and additional evaluation results, including complexity analysis, additional evaluation metrics and visual comparisons.
F.1 Training Details
F.1.1 Datasets
We train the model using three publicly available datasets: ImageNet-21K [40] contains 14.2M images across 21K categories. After removing images with a shorter edge below 256 pixels, 9.3M images remain for $256\times 256$ pre-training. OpenImages [27] and SA-1B [25] provide high-resolution images. We use 1.7M directly downloadable images from OpenImages V4 and all 11.1M images from SA-1B for both $256\times 256$ and $512\times 512$ resolution pre-training. Overall, CoD is trained on 22.1M images. Compared with modern diffusion pipelines, this is a relatively modest data scale, and the only filtering criterion is image resolution rather than extensive data cleaning. We expect that scaling up training with higher-quality data will further improve the performance of CoD.
F.1.2 Unified Training
Auxiliary Loss. Given the output of the condition decoder, we attach two lightweight auxiliary prediction heads to (1) reconstruct the input $x$ and (2) predict its DINO v2 [38] representation. The auxiliary objective is
@@ F.1.2 Unified Training
Figure 18: Ablation study on $\omega_{\alpha}$ in unified post-training.
Unified Post-Training. Motivated by the complementary strengths of unified training and auxiliary supervision, we combine both approaches to further improve performance. Specifically, we first pre-train CoD with auxiliary loss and then post-train it using unified training. For simplicity, rather than randomly selecting a subset of $\alpha\%$ samples for one-step training, we apply a unified objective to all samples during post-training: $\mathcal{L}_{\text{RF}}=\omega_{\alpha}\cdot\mathcal{L}_{\text{RF}}^{\text{one-step}}+\mathcal{L}_{\text{RF}}^{\text{multi-step}}$, where $\omega_{\alpha}$ serves as a weighting factor for the one-step loss component. This brings a little better performance. As shown in Table 5, this unified post-training strategy (ID = D) achieves the best overall results among all configurations.
We adopt $\omega_{\alpha}=1$ as the default setting for unified post-training. In Figure 18, we conduct a sensitivity analysis of this parameter, and the results demonstrate that a value of $\omega_{\alpha}=3$ yields additional performance gains, suggesting that further optimization of this hyperparameter could enhance overall efficacy.
Metric PerCo OSCAR DiffC DDCM OneDC CoD
User Score (1-5) $\uparrow$ 3.3 2.0 1.8 1.6 3.4 4.2
CLIP Sim. $\uparrow$ 0.84 0.78 0.72 0.63 0.84 0.89
Caption Sim. $\uparrow$ 0.87 0.83 0.76 0.72 0.86 0.89
VQA Acc. $\uparrow$ 69% 66% 58% 48% 71% 80%
Table 6: Additional evaluation metrics including user study and semantic scores around 0.004 bpp on Kodak.